OpenStudy (anonymous):

Find the cross product and show it is orthogonal to both u and v. u= <-3,2,3> v= <0,1,0> I dont understand! Can someone please help?

OpenStudy (anonymous):

Do you know how to find the cross product?

OpenStudy (anonymous):

\[ \langle-3,2,3\rangle \times \langle 0,1,0\rangle = \det\left( \begin{bmatrix} \mathbf i&\mathbf j&\mathbf k\\ -3&2&3\\ 0&1&0 \end{bmatrix} \right) \]

OpenStudy (anonymous):

\[u \times v=\]U x v=|■(i&j&k@-3&2&3@0&1&0)|={(2)(0)-(3)(1)}i-{(3)(0)-(-3)(0)}j+{(-3)(1)-(2)(0)}k =-3i-3k

OpenStudy (anonymous):

So the correct answer is -3i - 3k ?

OpenStudy (anonymous):

correct now again find dot product of (-3,2,3) and (-3,0,3) ,it should be (0,0,0) similarly find dot product of (0,1,0) with (-3,0,3),it should also be( 0,0,0)

OpenStudy (anonymous):

Okay great! Thanks for your help! I appreciate it

OpenStudy (anonymous):

Wait a minute, do you even know how to get the answer on your own?

OpenStudy (anonymous):

Yes I know how to get it on my own. He just explained it to me. Thanks for your concern

OpenStudy (anonymous):

correction it is <-3,0,-3> not (-3,0,3)

OpenStudy (anonymous):

If so, then what is the cross product of \(\langle 1,1,0\rangle\) and \(\langle 0,1,1\rangle\)?

OpenStudy (anonymous):

I can tell out right that isn't correct, because \(\langle 0,0,1\rangle \cdot \langle 0,1,1\rangle = 1\neq 0\)

OpenStudy (anonymous):

He showed me how to do it. That doesn't mean I am an expert at it. Sorry for the disappointment.

OpenStudy (anonymous):

Do you know the method? You don't need to be an expect.

OpenStudy (anonymous):

I certainly wouldn't understand how to do it based on his post alone. If you did then you're pretty clever.

OpenStudy (anonymous):

Find the cross product and show it is orthogonal to both u and v. u= 6k v= -i+3j+k This is the last question I have. I have attached the work below. I just want to make sure that I have it right.

OpenStudy (anonymous):

your cross product is correct but why are you finding unit normal to u cross v

OpenStudy (anonymous):

That is where I have my issue. I know how to do the steps, but I never know if my answer is write or not.

OpenStudy (anonymous):

again find dot product with u and v separately, if 0 then u cross v is orthogonal to both u and v.

OpenStudy (anonymous):

\[u.(u \times v)=\left( 6k \right).\left( -18i-6j \right)=0+0+0=0\] \[v.\left( u \times v \right)=\left( -i+3j+k \right).\left( -18i-6j \right)=\left( -1 \right)\left( -18 \right)+3\left( -6 \right)+1\left( 0 \right)=18-18=0\] i think now you are clear.

OpenStudy (anonymous):

When you find the dot product separately, you are looking for u and v to equal zero? Then, the answer would be 1?

OpenStudy (anonymous):

So, the cross product is 0? I am totally confused now :0

OpenStudy (anonymous):

I need to find the cross product, and show it is orthogonal to both u and v.

OpenStudy (anonymous):

you have found cross product as correct. Now you have to show this cross product is orthogonal to both i.e., u cross v is perpendicular to both u and v. for this you have to find dot product of u and u cross v. again find dot product of v with u cross v.

OpenStudy (anonymous):

this dot product as i have shown above should be 0.

OpenStudy (anonymous):